# A scalable adaptive deep Koopman predictive controller for real-time optimization of mixed traffic flow

### v1

@@ VI-B2 Computation time comparison
VI-C Real-time control optimization experiment for large-scale mixed traffic flow
@@
This section verifies the ability of AdapKoopPC for large-scale mixed traffic flow control, where the leading vehicle is followed by 50 vehicles. Three types of traffic flow scenarios covering different CAVs penetration rates, communication range degradation, and different CAVs distribution are set up to analyze the effectiveness and performance loss of AdapKoopPC in traffic flow optimization control in a wide range of scenarios.
VI-C1 CAVs penetration rates
@@
Fig. 12 depicts the overall evolution of the mixed vehicle with AdapKoopPC under different CAV penetration rates, including 0%, 10%, and 20%. Note that there are some white spacing lines, indicating that the trajectory here is a truck trajectory with a greater headway than the headway of the car. The following row shows the corresponding three-dimensional evolution diagram. The experimental comparison results can be mainly summarized into the following two points: a) When there is no CAV in the traffic flow, the disturbance of the leading vehicle propagates directly upstream with the traffic wave, which can be reflected by the increasingly prominent red area in Fig. 12(a). The three-dimensional velocity evolution diagram provides a more intuitive view of the poor driving conditions of the traffic flow. b) As the penetration rate increases to 10%, the traffic wave gradually dissipates and there is no significant fluctuation in the velocity of following vehicles. Until the penetration rate reaches 20%, the velocity oscillation can be almost ignored and maintained around the expected velocity of 25m/s. An experiment with the 30% penetration rate has also been conducted, and its effect is very close to that of 20%, which is not shown here. It can be considered that the effect is approaching saturation at 20% penetration rate. Overall, AdapKoopPC has the ability to significantly alleviate upstream disturbance waves.
VI-C2 Communication range degradation
@@
Another case has also been considered, where CAV deteriorates within communication range due to external factors. The vehicle can only receive information from the following vehicle. And the results are shown in Fig. 13. Note that more communication restricted scenarios fall between the communication degradation scenario discussed here and the communication lossless scenario, so the control effect of AdapKoopPC in other scenarios will be better than the extreme scenario presented below.
@@
Three penetration rates of 10%, 20%, and 30% are set for experiments under the premise of communication range degradation. Firstly, by comparing Fig. 13 (a) with Fig. 12 (a), AdapKoopPC still plays a significant role, and the traffic oscillation wave generated by the leading vehicle is significantly alleviated. Secondly, the difference between Fig. 13 (a) and Fig. 12 (b) intuitively reflects the impact of communication range degradation on the evolution improvement of the traffic flow. The oscillation slightly increases, indicating that the performance of AdapKoopPC has been slightly degraded. Similarly, in the scenario of a 20% penetration rate, CAV only considers optimizing and inducing the first vehicle behind it, resulting in a decrease in the overall mixed traffic flow evolution effect, which can be observed by comparing Fig. 13 (b) and Fig. 12 (c). Of course, under the above two penetration rates, the traffic flow oscillation propagation is still within an acceptable range, proving that AdapKoopPC can still induce and optimize the performance of the entire traffic flow even under the influence of communication range degradation. Fig. 13 (c) corresponds to a 30% penetration rate. The simulation results confirm that the evolution of the traffic flow is stable and smooth, and the optimization effect is approaching saturation.
VI-C3 CAVs distribution
@@
VI-C4 Number of controller deployments
We further investigate how many CAVs are required to deploy the proposed AdapKoopPC controller in random conditions to achieve a satisfactory level of traffic flow improvement. A mixed traffic flow consisting of 20 CAVs and 30 HDVs were randomly generated, where the AdapKoopPC was randomly deployed on the CAVs. In order to quantitatively analyze the traffic flow improvement, the velocity standard deviation and headway standard deviation of all vehicles in the mixed traffic flow are used as evaluation indicators. The results are shown in Table V. Overall, as the number of deployed controllers increases, the standard deviation of speed and headway gradually decreases. More importantly, the standard deviation of speed and headway archives convergence when the controller was 15 in such a mixed traffic flow. At such situation, the speed oscillation was reduced by 52.12%, and the standard deviation of the headway was reduced by 53.36%.
@@ VI-C4 Number of controller deployments
VII Conclusions
The main challenge currently faced by optimizing mixed traffic flow through CAVs is to accurately recognize and predict the driving behavior of HDVs while meeting real-time online computing capabilities. The Koopman operator theory and extended dynamic mode decomposition can achieve the linear approximation of nonlinear dynamic systems in high-dimensional space, providing feasible solutions for current issues. In this study, we propose AdapKoopnet to realistically describe and predict the HDVs driving behavior in complex scenarios. Furthermore, the AdaptKoopPC within a predictive control framework is developed to optimize real-time control of mixed traffic flow. AdapKoopnet has the capability to adaptively cluster potential driving scenarios and extract driving features from HDVs historical trajectories, which are then fused with the explicit state of the vehicle to form a high-dimensional representation of the state. The AdapKoopPC integrates AdapKoopnet and benefits from its linear characteristics, to achieve more efficient real-time control of CAVs in mixed traffic flow. Subsequently, trajectory prediction and traffic flow control optimization experiments are conducted, the following conclusions are mainly drawn:
Benefiting from the Koopman theory, the multi-step prediction performance of AdapKoopnet based on linear space surpasses LSTM-based nonlinear encoder-decoder of the same dimension. And the semantic extraction of trajectory context plays an indispensable positive role in the accuracy of HDVs state prediction;
With the appropriate architecture and loss function design of AdapKoopnet, it achieves powerful trajectory clustering and scenario recognition without pre-label, as well as adaptive driving characteristics extraction;
AdapKoopPC does not rely on prior car following model or pre-collected simulation trajectories, has better scenario generalization ability. In small-scale experiments, AdapKoopPC has shown promising results in slowing down traffic disturbance oscillations and improving traffic flow stability. Compared with baseline models, it has more satisfactory computational efficiency and practical application potential;
The large-scale experimental results show that when the penetration rate of CAVs is 20%, the inhibitory effect of AdapKoopPC on oscillations is close to saturation. Besides, in the case of communication degradation, the control effect of AdapKoopPC with 20% CAV penetration rate is roughly equivalent to that of a 10% CAV penetration rate without communication loss; Moreover, the control effect of AdaptKoopPC varies under different CAVs distributions, yet consistently exhibit superior performance in all cases.
There are some limitations of this study. For example, Koopman operator does not support online updates. And AdapKoopPC still faces dimensional disasters under ultra large scale mixed traffic flow. In future work, the online learning and efficient distributed control framework will be developed. Furthermore, expanding the application scenario of AdapKoopPC to multiple lanes is also worth further exploration.

### v3

@@ 6.2.2 Computation time comparison
6.3 Real-time control optimization experiment for large-scale mixed traffic system
@@
This section verifies the effectiveness and strong generalization ability of AdadKoopPC for large-scale mixed traffic system control, where the leading vehicle is followed by 50 vehicles. Three types of traffic flow scenarios covering different CAVs penetration rates, communication range degradation, and different CAVs distribution are set up to analyze the effectiveness and performance loss of AdadKoopPC in traffic flow optimization control in a wide range of scenarios.
6.3.1 CAVs penetration rates
@@
Fig. 12 depicts the overall evolution of the mixed vehicle with AdadKoopPC under different CAV penetration rates in the large-scale mixed traffic system scenario, including 0%, 10%, and 20%. The horizontal axis represents time, the vertical axis represents vehicle position, and the color axis represents velocity. Note that there are some white spacing lines, indicating that the trajectory here is a truck trajectory with a greater headway than the headway of the car. The following row shows the corresponding three-dimensional evolution diagram. The experimental comparison results can be mainly summarized into the following two points: a) When there is no CAV in the traffic system, the disturbance of the leading vehicle propagates directly upstream with the traffic wave, which can be reflected by the increasingly prominent red area in Fig. 12(a). The three-dimensional velocity evolution diagram provides a more intuitive view of the poor driving conditions of the traffic system. b) As the penetration rate increases to 10%, the traffic wave gradually dissipates and there is no significant fluctuation in the velocity of following vehicles. Until the penetration rate reaches 20%, the velocity oscillation can be almost ignored and maintained around the expected velocity of 25m/s. An experiment with the 30% penetration rate has also been conducted, and its effect is very close to that of 20%, which is not shown here. It can be considered that the effect is approaching saturation at 20% penetration rate. Overall, AdadKoopPC has the ability to significantly alleviate upstream disturbance waves.
6.3.2 Communication range degradation
@@
Another special case has also been considered, where CAV deteriorates within communication range due to external factors. The vehicle can only receive information from the following vehicle, and cannot obtain the driving states of all vehicles. And the performance degradation of AdadKoopPC in this scenario is investigated ,the results are shown in Fig. 13. Note that more communication restricted scenarios fall between the communication degradation scenario discussed here and the communication lossless scenario, so the control effect of AdadKoopPC in other scenarios will be better than the extreme scenario presented below.
@@
Three penetration rates of 10%, 20%, and 30% are set for experiments under the premise of communication range degradation. Firstly, by comparing Fig. 13 (a) with Fig. 12 (a), AdadKoopPC still plays a significant role, and the traffic oscillation wave generated by the leading vehicle is significantly alleviated. Secondly, the difference between Fig. 13 (a) and Fig. 12 (b) intuitively reflects the impact of communication range degradation on the overall evolution improvement of the traffic system. The oscillation slightly increases, indicating that the performance of AdadKoopPC has been slightly degraded. Similarly, in the scenario of a 20% penetration rate, due to the impact of communication range degradation on AdadKoopPC, CAV only considers optimizing and inducing the first vehicle behind it, resulting in a decrease in the overall mixed traffic system evolution effect, which can be observed by comparing Fig. 13 (b) and Fig. 12 (c). Of course, under the above two penetration rates, even if the communication range is degraded, the traffic system oscillation propagation is still within an acceptable range, proving that AdadKoopPC can still induce and optimize the performance of the entire system even under the influence of communication range degradation. Fig. 13 (c) corresponds to a 30% penetration rate. The simulation results confirm that the evolution of the traffic system is stable and smooth, and the optimization effect is approaching saturation.
6.3.3 CAVs distribution
@@
6.3.4 Number of controller deployments
We further investigate how many CAVs are required to deploy the proposed AdapKoopPC controller in random conditions to achieve a satisfactory level of traffic flow improvement. A mixed traffic flow consisting of 20 CAVs and 30 HDVs were randomly generated, where the AdapKoopPC was randomly deployed on the CAVs. In order to quantitatively analyze the traffic flow improvement, the velocity standard deviation and headway standard deviation of all vehicles in the mixed traffic flow are used as evaluation indicators. The results are shown in Table 5. Overall, as the number of deployed controllers increases, the standard deviation of speed and headway gradually decreases. More importantly, the standard deviation of speed and headway archives convergence when the controller was 15 in such a mixed traffic flow. At such situation, the speed oscillation was reduced by 52.12%, and the standard deviation of the headway was reduced by 53.36%.
@@ 6.3.4 Number of controller deployments
7 Conclusions
This study addressed the fundamental challenge of mitigating traffic oscillations in mixed traffic flow: the inherent conflict between the need for high-fidelity, nonlinear models of HDV behavior and the requirement for computationally tractable, linear models for real-time MPC. To resolve this, we proposed a novel framework based on Koopman operator theory, which successfully decouples modeling complexity from control complexity. Our proposed deep Koopman network, AdapKoopnet, demonstrated its capability to learn complex, personalized driving behaviors from naturalistic data, autonomously identifying latent traffic scenarios and driver characteristics without pre-defined labels. Its superior multi-step prediction performance, which surpasses that of traditional nonlinear deep learning models, provided the foundation for our control strategy.
The resulting AdapKoopPC framework integrates these learned, high-dimensional linear models into an MPC scheme, creating a scalable and adaptive control architecture for CAVs. Extensive simulations validated the efficacy of this approach. In both small- and large-scale mixed traffic systems, AdapKoopPC significantly attenuated traffic oscillations and enhanced traffic stability, proving effective even at low CAV penetration rates of 10-20%. The framework demonstrated robustness across various CAV distributions and under degraded communication conditions, all while maintaining remarkable computational efficiency suitable for real-time deployment. This work confirms that the proposed Koopman-based data-driven predictive control paradigm offers a powerful and practical solution for smoothing mixed traffic flow.
Despite these promising results, this study has limitations that open avenues for future research. The current AdapKoopnet relies on offline training and does not support online updates, which could be critical for adapting to non-stationary driving behaviors. Moreover, the scalability of the centralized MPC in AdapKoopPC may still face challenges in ultra-large-scale systems. Future work will focus on developing online learning mechanisms for the Koopman operator and exploring efficient distributed control architectures to further enhance the framework’s applicability to complex, real-world traffic environments, including multi-lane scenarios.
Acknowledgement
This research was supported by the project of the National Key R&D Program of China (No. 2023YFB4302701), the National Natural Science Foundation of China (No. 52232012, 52272343, 52131203), SEU Innovation Capability Enhancement Plan for Doctoral Students (No.CXJH_SEU 25178) and Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX24_0452).